OpenStudy (anonymous):

how do you evaluate this on your calculator log(4)23 but the 4 is lower than the word log

OpenStudy (anonymous):

\[\frac{ln(23)}{ln(4)}\]

OpenStudy (anonymous):

\[log_b(x)=\frac{log_a(x)}{log_a(b)}\]

OpenStudy (anonymous):

yeh remembering the formula is for people that dont know what they are doing though

OpenStudy (anonymous):

but you only have two logs on your calculator, \[\log=\log_{10}\] and \[\ln=\log_e\]

OpenStudy (anonymous):

i disagree. if i want to solve for example \[2^x = 1000\] for x the only way to do it is to compute \[\frac{\ln(1000)}{\ln(2)}\]

OpenStudy (anonymous):

because i have no way of knowing what \[\log_2(1000)\] is

OpenStudy (anonymous):

so without the change of base formula i am lost. it is how you solve an equation when the variable is in the exponent

OpenStudy (anonymous):

yeh but you derive the eqn, you dont memroise it

OpenStudy (anonymous):

maths is alot easy to learn derivations than to memorise eqns

OpenStudy (anonymous):

so how wuld i input this into my calc

OpenStudy (anonymous):

no i just know that \[A = b^x \] same as \[x=\frac{ln(A)}{\ln(x)}\]

OpenStudy (anonymous):

sorry i mean \[A=b^x \] means \[x=\frac{\ln(A)}{\ln(x)}\]

OpenStudy (anonymous):

lols :P

OpenStudy (anonymous):

damn \[A=b^x\] means \[x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

that is the one i want!

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

I just remember a basic proof A= b^x ln(A) = ln(b^x) ln(A) = x ln(b) x= ln(A) / ln(b)

OpenStudy (anonymous):

i do not need to reinvent the wheel every time and say "take the log, pull out the exponent, and then solve " because it is always the same

OpenStudy (anonymous):

takes 5seconds to derive, says me all the time smashing some formula into my head for no reason

OpenStudy (anonymous):

ok we will have to agree to disagree. same reason i would never say \[e^x = A\] \[\ln(e^x)=\ln(A)\]\[x\ln(e)=\ln(A)\] \[x=\ln(A)\]

OpenStudy (anonymous):

it is just how one solves for a variable in the exponent!

OpenStudy (anonymous):

the meaning of \[b^x\] is \[e^{x\ln(b)}\] elsewise \[2^{\sqrt{3}}\] makes no sense

OpenStudy (anonymous):

can someone atleast tell me waht the answer to the problem is

OpenStudy (anonymous):

yes it is the first thing i wrote \[x=\frac{\ln(23)}{\ln(4)}\]

OpenStudy (anonymous):

lols, whatever u wanna do. but thats partly the reason why so many people seems to struggle with maths, theythink that is all about remember formulas , and that there are so many formulas ! When all they really need is figure out what to remember and what not to remember lols . The more you can derive the less likely it is to make a mistake. Maths is just about smart memorisation lol

OpenStudy (anonymous):

the almighty "change of base" formula which is good to remember

OpenStudy (anonymous):

how would i put ln into calc

OpenStudy (anonymous):

depends on the calculator. in mine i would type \[ln (23) \div ln (4)\]

OpenStudy (anonymous):

make sure to close the parentheses about the 23

OpenStudy (anonymous):

is there an ln key

OpenStudy (anonymous):

yes there should be. what calculator are you using?

OpenStudy (anonymous):

you dont know how to use a calculator ? :|

OpenStudy (anonymous):

the teacher really dropped the ball somewhere lols

OpenStudy (anonymous):

im using ti-84, perhaps i just cant find the key

OpenStudy (anonymous):

fancy one!

OpenStudy (anonymous):

on the left hand side between log and sto

OpenStudy (anonymous):

neva mind i put on my glasses and found the key, thank u satellite for ur help w/o being rude

OpenStudy (anonymous):

no problem btw you could also type in \[log(23)\div log(4)\] and get the same answer

OpenStudy (anonymous):

i get 2.26 rounded

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

yay! gnight!

OpenStudy (anonymous):

u 2